# Quantum Statistical Effects on Warm Dark Matter and the Mass Constraint from the Cosmic Large Scale Structure

### v1

@@ Abstract
The suppression of small-scale matter power spectrum is a distinct feature of Warm Dark Matter (WDM), which permits a constraint on the WDM mass from galaxy surveys. In the thermal relic WDM scenario, quantum statistical effects are not manifest. In a unified framework, we investigate the quantum statistical effects for a fermion case with a degenerate pressure and a boson case with a Bose-Einstein condensation (BEC). Compared to the thermal relic case, the degenerate fermion case only slightly lowers the mass bound while the boson case with a high initial BEC fraction ($\gtrsim 90\%$) significantly lowers it. On the other hand, the BEC fraction drops during the relativistic-to-nonrelativistic transition and completely disappears if the initial fraction is below $\sim 64$%. Given the rising interest in resolving the late-time galaxy-scale problems with boson condensation, a question is posed on how a high initial BEC fraction can be dynamically created so that a DM condensed component remains today.
@@ 1 Introduction
A wide range of cosmological and astronomical phenomena from galaxy scales (Zwicky, 1933; Smith, 1936; van de Hulst et al., 1957; Clowe et al., 2006; Lin & Ishak, 2016) to large scales (Dark Energy Survey Collaboration et al., 2018; Hikage et al., 2019; Joudaki et al., 2018; Ade & others (2020), Planck Collaboration) point to the existence of dark matter (DM). However, little is known about dark matter except that they are around five times in mass as baryons and are highly nonrelativistic at the time of recombination (i.e., cold). Here, we assume that dark matter is some unknown particles.
@@
An important parameter of DM is the particle mass, which is related to almost all aspects of the DM study such as model buildings, early-universe production mechanisms, and direct and indirect searches. The cosmic large-scale structure (LSS) has been a powerful probe of DM properties, especially the “warmness” of DM. The DM early-time thermal motion smooths out the structures and prevents gravitational growth at small scales. In the thermal relic scenario, WDM became a free stream when they were still relativistic. Matching the current DM energy density requires that lighter DM be warmer. So, the nondetection of the suppression of the matter power spectrum puts a lower bound of the WDM mass (Colombi et al., 1996), which is currently of a few keV (Iršič et al., 2024).
@@
Such a constraint is not absolute but depends on the assumption of the DM production mechanism (Bhupal Dev et al., 2014). In particular, while a fermion WDM is assumed in the usual thermal relic scenario, quantum statistical effects only play a minor role. When the WDM number density is larger than that in the thermal relic scenario, quantum statistical effects begin to be significant. For the fermion case, a degenerate pressure is added on top of the thermal pressure (Bar et al., 2021; Carena et al., 2022). For the boson case, a Bose-Einstein condensation (BEC) component appears (Madsen, 1992) and can affect the cosmic background evolution and structure formation (Fukuyama & Morikawa, 2006; Harko, 2011; Chavanis, 2012).
@@
Given the importance of the WDM mass constraint and the growing interest in quantum effects on galaxy and cosmological scales, in this work, we investigate how quantum statistical effects impact the mass constraint from LSS. From the view of conservation of specific entropy, we analyze the evolution of quantum statistical effects during the relativistic-to-nonrelativistic transition (RNRT) and their impacts on RNRT and LSS. In particular, compared previous works on cosmological DM condensation (Fukuyama & Morikawa, 2006; Harko, 2011), we focus on how the BEC fraction evolves during RNRT and the effects on the WDM mass constraint from LSS. The scenario considered in this work serves as a minimum extension to the thermal relic WDM that considers both a degenerate fermion case and a boson case with a BEC component. We dub it the quantum Warm Dark Matter (qWDM).
@@ 2 Analysis
with a mass $m$, intrinsic degree of freedom $g$, temperature $T$, chemical potential $\mu$. In the denominator, $+1$ is for fermion and $-1$ for boson. Note that this DM temperature $T$ is different from the standard-sector temperature. In this work, we do not assume any early thermal contact between DM and the standard model (SM) and thus DM remains hidden. [^1] It is more suitable to call our scenario the “Hidden Dark Matter” (Chen & Tye, 2006), which is more generalized. We adopted the term WDM to better highlight the quantum statistical effects compared to the thermal relic scenario. Since we assume DM has an equilibrium phase-space distribution instead of being a free stream, it requires some DM-DM interaction to keep DM in equilibrium at least in the early time (Egana-Ugrinovic et al., 2021). We take $g=2$ for the fermion case and $g=1$ for the boson case, but the conclusions obtained in this work can be readily generalized to cases with a higher $g$.
We define the following two dimensionless variables,
@@
Then, $x$ is the mass-to-temperature ratio and $\Delta$ is the effective chemical potential-to-temperature. We call $\mu-m$ the effective chemical potential for the following reason. In the relativistic limit, the mass $m$ can be ignored. In the nonrelativistic limit, the distribution Eq. (1) reduces to the nonrelativistic form, from which we can identify $\mu-m$ as the nonrelativistic chemical potential.
From Eq. (1), we calculate the physical particle number density $n$, energy density $\rho$, pressure $P$ as well as entropy density $s$ for given temperature and chemical potential; see appendix A. Some important results for $n$, $\rho$, $P$, and $s/n$ in the relativistic and nonrelativistic limits are summarized in Table 1. Those limits are useful for obtaining the relations between the initial and final values of dynamical variables and verifying our numerical solutions of the qWDM background evolution.
@@
where $n_{\textsc{bec}}$ is the particle number density of qWDM in the BEC state and $n_{\rm tot}=n+n_{\textsc{bec}}$ is the total qWDM number density. We assume that there is no internal energy state and the BEC component of DM condenses into the zero-momentum state so that the total energy density is
@@ 2.1 The DM energy density today
$\alpha\equiv\frac{T_{0}}{T_{\gamma}^{0}}$ with $T_{0}\equiv\lim_{a\rightarrow 0}aT$ and $T_{\gamma}^{0}$ is the photon temperature today. Please see Table 1 for the definition of the $J$’s functions. Now, the superscript “--” is for fermion and “$+$” for boson. Note that, while we matched the form of Eq. (8) to that of Eq. (6) in Colombi et al. (1996), the $\beta$ here is determined by the initial effective chemical potential-to-temperature ratio and the initial fraction of BEC component instead of an arbitrary factor in the momentum distribution. For the same initial temperature, Both a higher degree of degeneracy for the fermion case and a higher BEC fraction for the boson case give a higher $\beta$ factor and a higher DM density today. However, $\alpha$ is not directly related to an observable and we shall replace it in Eq. (8) with the RNRT scale factor, which predominantly determines the suppression scale of the matter power spectrum.
@@ 2.2 The relativistic-to-nonrelativistic transition
In general, qWDM starts with a relativistic phase where $c_{\rm s}$ equals $1/\sqrt{3}$ and then transitions to a nonrelativistic phase where it drops as $1/a$. We define the transition scale factor $a_{\rm nr}$ by the following asymptotic behavior of $c_{\rm s}$ in the nonrelativistic limit,
@@
Different quantum statistics have distinct effects on $a_{\rm nr}$. The fiducial case corresponding to (but somewhat different from) the thermal relic WDM scenario is a fermion case with a vanishing initial chemical potential and $g=2$. This case has only a small difference from the classical case on the value of $a_{\rm nr}$ for the same initial conditions (Lin et al., 2023). We show the evolution of $c_{\rm s}$ for the classical case by the orange curve in Figure 1.
Figure 1: Effects from quantum statistics on the adiabatic sound speed. The degenerate pressure postpones the relativistic-to-nonrelativistic transition for the fermion case, while a BEC component for the boson case expedites it. The same initial value of $x/a$ is applied.
For the fermion case with a positive initial $\Delta$ (denoted as $\Delta_{\rm i}$), the degenerate pressure postpones RNRT with a larger $a_{\rm nr}$ compared to the classical case. The larger $\Delta_{\rm i}$, the larger $a_{\rm nr}$. The reason is the following. In the presence of the degenerate pressure, the condition $m\gg T$ (or $x\gg 1$) is no longer sufficient to define the nonrelativistic phase as it would be in the classical case. If the degenerate pressure is high enough, a large fraction of DM particles can still occupy relativistic momentum states. Mathematically, only when $x\gg\Delta$ will the momenta of most particles be smaller than the mass and Eq. (1) be approximated by a nonrelativistic form. Therefore, the nonrelativistic limit requires both $x\gg 1$ and $x\gg\Delta$, that is, in addition, the particle mass needs to be much larger than the effective chemical potential. During the relativistic phase, $x$ increases linearly with $a$, but $\Delta$ remains a constant. Only when $x$ increases to be larger than $\Delta$ will DM enter the nonrelativistic phase. As a result, the higher the initial degree of degeneracy (i.e., the larger $\Delta_{\rm i}$), the later the transition. Such an effect is shown by the blue curve in Figure 1.
For the boson case with a fraction of the BEC component, an opposite effect takes place and the transition is advanced compared to the classical case. This is because now $x\ll 1$ is not the criteria for the relativistic phase. When $x\ll 1$ with the presence of a BEC component, the total energy density is,
@@
where both $J_{3}^{+}(0)$ and $J_{4}^{+}(0)$ are of the order of unity. Meanwhile, $P=\rho/3$ is still satisfied for $x\ll 1$. When $x\sim(1-r)/r$, the term $mn_{\textsc{bec}}$ is comparable to $\rho$ and the pressure. This estimate breaks down at small $r$ values where higher order terms in $x$ are needed. Numerical analyses suggest that such an estimate is good for $r\sim 1$ and that $x\ll(1-r)$ is a good criterion for the relativistic phase for all $r>0$. Then, once $x$ increases to $(1-r)$, qWDM already begins to enter the nonrelativistic phase. Therefore, the larger $r_{\rm i}$, the earlier the RNRT.
For all cases, the transition scale is proportional to the initial value of $a/x$, that is,
@@
The coefficient $C_{\rm cs}$ depends on the initial degree of degeneracy for the fermion case and the initial BEC fraction for the boson case. It can be derived analytically. This is done by first substituting the nonrelativistic forms of $\rho$ and $P$ (see Table 1) into Eq. (10) to relate $c_{\rm s}$ and $x$ in the nonrelativistic phase. Then, the asymptotic evolution of $x$ at the nonrelativistic phase can be derived by matching the initial and the final comoving particle density. As a result, it can be shown that,
@@
The larger $C_{\rm cs}$, the later the transition. We show $C_{\rm cs}$ as function of $\Delta_{\rm i}$ and $r_{\rm i}$ in Figure 2.
Figure 2: The effect of quantum statistics on the coefficient $C_{\rm cs}$ defined in Eq. (13). The larger $C_{\rm cs}$, the later the relativistic-to-nonrelativistic transition for a given initial mass-to-temperature ratio.
By combining Eqs. (8) and (13), we relate the DM energy fraction to the transition scale factor which reads
@@
where $C_{\rm cs}^{0}$ and $\Delta_{\rm f}^{0}$ are the values of $C_{\rm cs}$ and $\Delta_{\rm f}$ in the fiducial case (i.e., a fermion case with $\Delta_{\rm i}=0$). The value of $\hat{a}_{\rm nr}$ is motivated by the discussion on the suppression scale in the next section.
2.3 The suppression scale of matter power spectrum
@@
We find that $A=\hat{a}_{\rm nr}=1.8\times 10^{-7}$ and $B=0.16$ make Eq. (19) fit the numerical results well. [^3] Besides $a_{\rm nr}$ that dominantly determines the suppression scale, different quantum statistics also make the transition width somewhat different; see Lin et al. (2023). While we ignored such a small difference, including it does not qualitatively change our conclusions. For a given qWDM mass, $a_{\rm nr}$ can be inferred form Eq. (15), which is put in Eq. (19) to give $\ell_{\rm s}$. Despite the different physical settings, the $\ell_{\rm s}$ inferred for a given mass with $\Delta_{\rm i}=0$ and $g=2$ is only a factor of $\sim 1.5$ different compared to that in the free-streaming thermal relic WDM scenario provided in other works such as Bode et al. (2001); Hansen et al. (2002); Viel et al. (2013).
@@
$K(y)=y\,(1-0.16\ln y)\,.$ (21)
Importantly, since LSS constrains $\ell_{\rm s}$, from Eq. (20) we can see that the quantum statistical effects on the constraint of the WDM mass are reflected in $\gamma$. Then, from Eq. (16), there are the two dominant factors that affect $\gamma$ and hence the mass constraint: (1) $\beta$, that is the ratio of number density between the qWDM case and the thermal relic WDM case; (2) $C_{\rm cs}$, that determines the time of RNRT.

### v2

# Quantum Statistical Effects on Warm Dark Matter and the Mass Constraint from the Cosmic Structure at Small Scales

@@ Abstract
The suppression of the small-scale matter power spectrum is a distinct feature of Warm Dark Matter (WDM), which permits a constraint on the WDM mass from galaxy surveys. In the thermal relic WDM scenario, quantum statistical effects are not manifest. In a unified framework, we investigate the quantum statistical effects for a fermion case with a degenerate pressure and a boson case with a Bose-Einstein condensation (BEC). Compared to the thermal relic case, the degenerate fermion case only slightly lowers the mass bound, while the boson case with a high initial BEC fraction ($\gtrsim 90\%$) significantly lowers it. On the other hand, the BEC fraction drops during the relativistic-to-nonrelativistic transition and completely disappears if the initial fraction is below $\sim 64$%. Given the rising interest in resolving the late-time galaxy-scale problems with boson condensation, a question is posed on how a high initial BEC fraction can be dynamically created so that a condensed DM component remains today.
@@ 1 Introduction
A wide range of cosmological and astronomical phenomena from galaxy scales (Zwicky, 1933; Smith, 1936; van de Hulst et al., 1957; Clowe et al., 2006; Lin & Ishak, 2016; Salucci, 2019) to large scales (Abbott et al., 2018; Hikage et al., 2019; Joudaki et al., 2018; Ade & others (2020), Planck Collaboration) point to the existence of dark matter (DM). However, little is known about DM, except that it is around five times the mass of baryons and highly nonrelativistic at the time of recombination (i.e., cold). Here, we assume that dark matter is some unknown particle.
@@
An important parameter of DM is the particle mass, which is related to almost all aspects of the DM study, such as model building, early-universe production mechanisms, and direct and indirect searches. The cosmic structure, especially at intermediate to small scales, has been a powerful probe of DM properties, especially the “warmness” of DM. The early-time thermal motion of DM smooths out the structures and prevents gravitational growth at small scales. In the standard thermally produced relic scenario, WDM became a free stream when it was still relativistic. Matching the current DM energy density requires that lighter DM be warmer. So, the nondetection of the suppression of the matter power spectrum places a lower bound on the WDM mass (Colombi et al., 1996), which is currently of a few keV from the Lyman-$\alpha$ forest (Iršič et al., 2024; Villasenor et al., 2023). While we focus on the Lyman-$\alpha$ constraint, we remind readers that the suppression of the matter power spectrum, and thus the WDM mass, can also be constrained by other observations. These include the abundance of small halos and subhalos as probed by strong lensing (Li et al., 2016), subhalo-star stream interactions (Erkal & Belokurov, 2015a, b), satellite galaxy abundance (Nadler et al., 2021; Newton et al., 2020), the high-redshift luminosity function (Rudakovskyi et al., 2021), and, potentially, the global and anisotropic 21cm signals (Boyarsky et al., 2019; Chatterjee et al., 2019).
@@
The above constraint on the DM mass is not absolute but depends on the assumption of the DM production mechanism (Bhupal Dev et al., 2014). In particular, while a fermion WDM is assumed in the usual thermal relic scenario, quantum statistical effects only play a minor role. When the WDM number density is larger than that in the thermal relic scenario, quantum statistical effects begin to be significant. For the fermion case, a degenerate pressure is added on top of the thermal pressure (Bar et al., 2021; Carena et al., 2022). For the boson case, a Bose-Einstein condensation (BEC) component appears (Madsen, 1992) and can affect the cosmic background evolution and structure formation (Fukuyama & Morikawa, 2006; Harko, 2011; Chavanis, 2012).
@@
Given the importance of the WDM mass constraint and the growing interest in quantum effects on galaxy and cosmological scales, in this work, we investigate how quantum statistical effects impact the mass constraint from the Lyman-$\alpha$ forest. From the view of conservation of specific entropy, we analyze the evolution of quantum statistical effects during the relativistic-to-nonrelativistic transition (RNRT) and their impacts on RNRT and the matter power spectrum. In particular, compared to previous works on cosmological DM condensation (Fukuyama & Morikawa, 2006; Harko, 2011), we focus on how the BEC fraction evolves during RNRT and the effects on the WDM mass constraint from cosmic structure. The scenario considered in this work serves as a minimum extension to the standard thermal relic WDM that considers both a degenerate fermion case and a boson case with a BEC component. We dub it quantum WDM (qWDM).
@@ 2 Analysis
with a mass $m$, intrinsic degree of freedom $g$, temperature $T$ and chemical potential $\mu$. In the denominator, $+1$ is for fermion and $-1$ for boson. Note that this DM temperature $T$ is different from the standard-sector temperature. In this work, we do not assume any early thermal contact between DM and the standard model, and thus DM remains hidden. [^1] It is more suitable to call our scenario the “Hidden Dark Matter” (Chen & Tye, 2006), which is more generalized. We adopted the term WDM to better highlight the quantum statistical effects compared to the thermal relic scenario. Since we assume that DM has an equilibrium phase-space distribution instead of being a free stream, it requires some DM-DM interaction to keep DM in equilibrium at least in the early time (Egana-Ugrinovic et al., 2021). We take $g=2$ for the fermion case and $g=1$ for the boson case, but the conclusions obtained in this work can be readily generalized to cases with a higher $g$.
We define the following two dimensionless variables:
@@
Then $x$ is the mass-to-temperature ratio, and $\Delta$ is the effective chemical potential-to-temperature ratio. We call $\mu-m$ the effective chemical potential for the following reason. In the relativistic limit, the mass $m$ can be ignored. In the nonrelativistic limit, the distribution Eq. (1) reduces to the nonrelativistic form, from which we can identify $\mu-m$ as the nonrelativistic chemical potential.
From Eq. (1), we calculate the physical particle number density $n$, energy density $\rho$, and the pressure $P$, as well as the entropy density $s$ for given temperature and chemical potential; see Appendix A. Some important results for $n$, $\rho$, $P$, and $s/n$ in the relativistic and nonrelativistic limits are summarized in Table 1. Those limits are useful for obtaining the relations between the initial and final values of dynamical variables and verifying our numerical solutions of the qWDM background evolution.
@@
where $n_{\textsc{bec}}$ is the particle number density of qWDM in the BEC state and $n_{\rm tot}=n+n_{\textsc{bec}}$ is the total qWDM number density. We assume that there is no internal energy state, and the BEC component of DM condenses into the zero-momentum state so that the total energy density is
@@ 2.1 The DM energy density today
$\alpha\equiv\frac{T_{0}}{T_{\gamma}^{0}}$ with $T_{0}\equiv\lim_{a\rightarrow 0}aT$ and $T_{\gamma}^{0}$ is the photon temperature today and
$J_{s}^{\mp}(\Delta)\equiv\int_{0}^{\infty}\frac{z^{s-1}{\rm d}z}{\exp(z-\Delta)\pm 1}\,.$ (10)
Now, the superscript “--” is for fermion and “$+$” for boson. Note that, while we matched the form of Eq. (8) to that of Eq. (6) in Colombi et al. (1996), the $\beta$ here is determined by the initial effective chemical potential-to-temperature ratio and the initial fraction of BEC component instead of an arbitrary factor in the momentum distribution. Unlike in the free-streaming case, the $f(p)$ in our scenario follows an explicit nonrelativistic equilibrium form. Consequently, $T_{0}$ differs from the current temperature of qWDM. For the same initial temperature, both a higher degree of degeneracy for the fermion case and a higher BEC fraction for the boson case give a higher $\beta$ factor and a higher DM density today. However, $\alpha$ is not directly related to an observable, and we shall replace it with the RNRT scale factor, which predominantly determines the suppression scale of the matter power spectrum.
@@ 2.2 The relativistic-to-nonrelativistic transition
In general, qWDM starts with a relativistic phase where $c_{\rm s}$ equals $1/\sqrt{3}$ and then transitions to a nonrelativistic phase where it drops as $1/a$. We define the transition scale factor $a_{\rm nr}$ by the following asymptotic behavior of $c_{\rm s}$ in the nonrelativistic limit:
@@
Different quantum statistics have distinct effects on $a_{\rm nr}$. The fiducial case corresponding to (but somewhat different from) the thermal relic WDM scenario is a fermion case with a vanishing initial chemical potential and $g=2$. This case has only a small difference from the classical case on the value of $a_{\rm nr}$ for the same initial conditions (Lin et al., 2023). We show the evolution of $c_{\rm s}$ for the classical case with the orange curve in Figure 1.
Figure 1: Effects of quantum statistics on the adiabatic sound speed. The degenerate pressure postpones the RNRT for the fermion case, while a BEC component for the boson case expedites it. The same initial value of $x/a$ is applied.
For the fermion case with a positive initial $\Delta$ (denoted as $\Delta_{\rm i}$), the degenerate pressure postpones RNRT with a larger $a_{\rm nr}$ compared to the classical case. The larger $\Delta_{\rm i}$, the larger $a_{\rm nr}$. The reason is the following. In the presence of the degenerate pressure, the condition $m\gg T$ (or $x\gg 1$) is no longer sufficient to define the nonrelativistic phase, as it would be in the classical case. If the degenerate pressure is high enough, a large fraction of DM particles can still occupy relativistic momentum states. Mathematically, only when $x\gg\Delta$ will the momenta of most particles be smaller than the mass and Eq. (1) be approximated by a nonrelativistic form. Therefore, the nonrelativistic limit requires both $x\gg 1$ and $x\gg\Delta$; that is, in addition, the particle mass needs to be much larger than the effective chemical potential. During the relativistic phase, $x$ increases linearly with $a$, but $\Delta$ remains a constant. Only when $x$ increases to be larger than $\Delta$ will DM enter the nonrelativistic phase. As a result, the higher the initial degree of degeneracy (i.e., the larger $\Delta_{\rm i}$), the later the transition. Such an effect is shown by the blue curve in Figure 1.
For the boson case with a fraction of the BEC component, the opposite effect takes place, and the transition is advanced compared to the classical case. This is because now $x\ll 1$ is not the criteria for the relativistic phase. When $x\ll 1$ with the presence of a BEC component, the total energy density is,
@@
where both $J_{3}^{+}(0)$ and $J_{4}^{+}(0)$ are of the order of unity. Meanwhile, $P=\rho/3$ is still satisfied for $x\ll 1$. From Eq. (13), we can see that when $x\sim(1-r)/r$, the term $mn_{\textsc{bec}}$ is comparable to $\rho$ and the pressure. This estimate breaks down at small $r$ values where higher-order terms in $x$ are needed. Numerical analyses suggest that such an estimate is valid for $r\sim 1$ and that $x\ll(1-r)$ is a better criterion for the relativistic phase for all $r>0$. Then, once $x$ increases to $(1-r)$, qWDM already begins to enter the nonrelativistic phase. Therefore, the larger the $r_{\rm i}$, the earlier the RNRT.
For all cases, the transition scale is proportional to the initial warmness, that is,
@@
The coefficient $C_{\rm cs}$ depends on the initial degree of degeneracy for the fermion case and the initial BEC fraction for the boson case. It can be derived analytically. This is done first by substituting the nonrelativistic forms of $\rho$ and $P$ (see Table 1) into Eq. (11) to relate $c_{\rm s}$ and $x$ in the nonrelativistic phase. Then, the asymptotic evolution of $x$ at the nonrelativistic phase can be derived by matching the initial and the final comoving particle density. As a result, it can be shown that
@@
The larger the $C_{\rm cs}$, the later the transition. We show $C_{\rm cs}$ as function of $\Delta_{\rm i}$ and $r_{\rm i}$ in Figure 2.
Figure 2: The effect of quantum statistics on the coefficient $C_{\rm cs}$ defined in Eq. (14). The larger the $C_{\rm cs}$, the later the RNRT for a given initial mass-to-temperature ratio.
By combining Eqs. (8) and (14), we relate the DM energy fraction to the transition scale factor, which reads
@@
where $C_{\rm cs}^{0}$ and $\Delta_{\rm f}^{0}$ are the values of $C_{\rm cs}$ and $\Delta_{\rm f}$ in the fiducial case (i.e., a fermion case with $\Delta_{\rm i}=0$). The value of $\hat{a}_{\rm nr}$ is motivated by the discussion of the suppression scale in the next section.
2.3 The suppression scale of the matter power spectrum
@@
We find that $A=\hat{a}_{\rm nr}=1.5\times 10^{-7}$ and $B=0.145$ make Eq. (20) fit the numerical results well. [^3] Besides $a_{\rm nr}$, which predominantly determines the suppression scale, different quantum statistics also make the transition width somewhat different; see Lin et al. (2023). While we ignored such a small difference, including it does not qualitatively change our conclusions. For a given qWDM mass, $a_{\rm nr}$ can be inferred from Eq. (16), which is put in Eq. (20) to give $\ell_{\rm s}$. Despite the different physical settings, the $\ell_{\rm s}$ inferred for a given mass with $\Delta_{\rm i}=0$ and $g=2$ is only a factor of $\sim 1.5$ different compared to that in the free-streaming thermal relic WDM scenario provided in other works, such as Bode et al. (2001); Hansen et al. (2002); Viel et al. (2013). Figure 3 shows the ratios of the matter power spectrum between the qWDM and CDM cases, both with a DM mass of $m=5$ keV. The cutoff shape for different qWDM scenarios is similar to that of the standard thermal relic WDM. At the smallest scales, the matter power spectrum for qWDM exhibits oscillations. While this feature is not currently relevant for observations, it may become significant as smaller scales are explored in the future.
Figure 3: The ratio of matter power spectrum between some selected qWDM cases and the CDM case (solid curves). We set $m=5$ keV for all cases. The black-dashed curve corresponds to the standard thermal relic WDM with the same mass. The dotted curves are approximations obtained with Eqs. (20), (23) and (24). The red dotted curve is obtained by the same procedure but with the DM mass scaled by a factor of $\frac{1}{0.84}$; see Eq. (25).
@@
$K(y)=y\,(1-0.145\ln y)\,.$ (22)
Importantly, since the Lyman-$\alpha$ forest constrains $\ell_{\rm s}$, from Eq. (21) we can see that the quantum statistical effects on the constraint of the WDM mass are reflected in $\gamma$. Then, from Eq. (17), there are the two dominant factors that affect $\gamma$ and hence the mass constraint: (1) $\beta$, which is the ratio of the number density between the qWDM case and the thermal relic WDM case, and (2) $C_{\rm cs}$, which determines the time of RNRT.